OpenStudy (anonymous):

how do I simplify (4-sqrt(80))/4

OpenStudy (anonymous):

do you know what is \[\sqrt{80}\] in simplified form

OpenStudy (studygurl14):

\(\Large\frac{4-\sqrt{80}}{4}\rightarrow\frac{4-\sqrt{16\times5}}{4}\rightarrow???\) Can you do the rest @robyncurry7 ?

OpenStudy (anonymous):

\[4-16\sqrt{5}/4?\]

OpenStudy (studygurl14):

not exactly. \(\large\sqrt{16}=4\) So... \(\Large\frac{4-4\sqrt{5}}{4}\)

OpenStudy (studygurl14):

What next?

OpenStudy (studygurl14):

@robyncurry7